Solve: 9-3(2x-5)=4x-6 I got the answer 3 but I don't know why.
Please show me step by step to get the answer. Thanks

Answer by pedjajov(51)

You have the correct answer. Here is how to get it:
9-3(2x-5)=4x-6 , first do the multiplication to get rid of parentheses
, here you should know that negative number -3
multiplied with 2x gives -6x
multiplied with -5 gives +10
negative times negative is positive
positive times positive is positive
negative times positive is negative
positive times negative is negative
9-6x+15 =4x-6 , now combine 9 and 15 on the left side
24-6x =4x-6 , we can exchange left and right sides,
4x-6 =24-6x , add 6x to the both sides
4x-6+6x =24-6x+6x , combine terms with x on both sides
10x-6 =24 , add 6 to the both sides
10x-6+6 =24+6 , combine numbers on both sides
10x =30 , divide both sides by 10
10x/10 =30/10 , do division
x =3 , here is your answer :-)